OpenStudy (anonymous):

MEDALLLLLL!!!!!!! if h(x)=2^x then find x when h(x)=0.25

OpenStudy (okdutchman7):

You just have to put 0.25 in for x Which would be 2^0.25

OpenStudy (anonymous):

How

OpenStudy (anonymous):

#okdutchman7

OpenStudy (anonymous):

@okdutchman7

OpenStudy (anonymous):

h(x) is a function not x

OpenStudy (okdutchman7):

okay, are you in algebra 2?

OpenStudy (anonymous):

yes

OpenStudy (okdutchman7):

Have you studied logarithms yet?

OpenStudy (anonymous):

no

OpenStudy (okdutchman7):

Okay

OpenStudy (okdutchman7):

You have h of x which is a function. This function is equal to 2^x

OpenStudy (okdutchman7):

You have to flip both sides of the equation, since you are dealing with negative exponents.

OpenStudy (okdutchman7):

Can you tell me what x is in this equation? 2^x = 4

OpenStudy (anonymous):

its a power so x is 2

OpenStudy (okdutchman7):

Correct, but since the exponents are negative you have to flip it.

OpenStudy (anonymous):

so....

OpenStudy (okdutchman7):

instead of 2^2 = 4 the equation can be flipped to equal 2^-2 = 1/4

OpenStudy (okdutchman7):

Your answer would be -2

OpenStudy (okdutchman7):

Do you understand now?

OpenStudy (okdutchman7):

I hope I helped. : )

OpenStudy (anonymous):

thnks